# AHA! Animating Human Avatars in Diverse Scenes with Gaussian Splatting

### v1

@@ 3.1 Gaussian Reconstruction
Step 1: Canonical Gaussian parameterization on SMPL. Given multi-view images of a person performing diverse poses, we learn a mapping from SMPL poses to 3D Gaussians in posed space. Following Li et al. (2023b), we place Gaussians on the 2D manifold of the SMPL surface by constructing an approximate front–back UV atlas via orthographic projections of the SMPL mesh. Let $\bm{\beta}$ denote SMPL shape and $\bm{\theta}_{t}$ the pose at time $t$. We rasterize pose-conditioned features into a pose map $P_{t}\in\mathbb{R}^{H\times W\times C}$ - denoted using $\mathcal{M}(\bm{\beta},\bm{\theta_{t}})$. A per-identity StyleUNet $f_{\phi}$ predicts a set of canonical human Gaussians anchored on the SMPL surface: $\mathcal{G}^{\mathsf{C}}_{t}\;=\;f_{\phi}(P_{t})\;=\;\big\{(\mathbf{x}_{k}^{\mathsf{C}},\bm{\Sigma}_{k}^{\mathsf{C}},\mathbf{c}_{k},\alpha_{k})\big\}_{k=1}^{N_{H}}.$
@@
Step 3: Pose-to-Gaussian inference (test-time). Following Li et al. (2023b), we compute the top $K\!\in\![10,20]$ PCA components of training pose maps $\{P_{t}\}$, yielding mean $\bar{P}$ and basis $Q_{K}$. At inference, poses synthesized by our Gaussian-aligned motion module (Sec. 3.2) are first projected to this subspace and then mapped to posed Gaussians: $\tilde{P}_{y}\;=\;\bar{P}+Q_{K}z_{y}\qquad\mathcal{G}^{\mathsf{P}}_{y}\;=\;\mathrm{LBS}_{\bm{\theta}_{y}}\!\big(f_{\phi}(\pi\!\big(\mathcal{M}(\beta,\bm{\theta}_{y})\big))\big).$ Here we use $\pi$ to denote the projection to the subspace operation and $y$ to indicate a test-time pose. For further details please see supplementary materials.
@@ 3.2 Gaussian-Aligned Motion Synthesis
Design overview. We reuse a strong latent motion backbone trained on large scale mocap dataset and add two 3DGS-specific controllers: an RL locomotion policy that navigates between waypoints while avoiding scene Gaussians, and a deterministic latent-space optimizer that executes short, fine-grained actions near targets (e.g., stop, sit, grasp) before returning control to RL. While this explicit decomposition is not typical in existing motion synthesis frameworks, we find it especially effective in 3DGS settings, as this allows us to exploit the fact that in 3DGS scenes much of the raw scene detail can be abstracted to (i) a set of paths for navigation and (ii) action points (e.g., sitting locations, grasping targets provided by an animator) at which specific behaviors are executed - thus allowing for scene-aware motion synthesis without human-scene paired data. Both submodules operate consistently in the latent space of a learned motion model Zhao et al. (2025).
@@
Latent motion backbone. We adopt a latent motion prior, following prior work Zhao et al. (2025) trained on AMASS Punnakkal et al. (2021); Mahmood et al. (2019). Specifically, the model learns a compact motion-primitive space with a transformer VAE trained on mocap data, and places a diffusion prior in this latent space. Given motion history $\mathbf{H}$ and a future motion segment $\mathbf{X}$, the encoder $\mathcal{E}$ outputs a Gaussian posterior $q_{\phi}(\mathbf{z}\mid\mathbf{H},\mathbf{X})=\mathcal{N}(\bm{\mu},\sigma^{2}\mathbf{I})$ with reparameterized sample $\mathbf{z}=\bm{\mu}+\sigma\odot\bm{\varepsilon}$ where $\bm{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The decoder $\mathcal{D}$ reconstructs motion as $\widehat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\mathbf{z})$. On this latent space, a denoiser $\mathcal{G}$ operates with forward process $q(\mathbf{z}_{\tau}\mid\mathbf{z}_{\tau-1})=\mathcal{N}(\sqrt{1-\beta_{\tau}}\,\mathbf{z}_{\tau-1},\,\beta_{\tau}\mathbf{I})$ and predicts the clean code $\widehat{\mathbf{z}}_{0}=\mathcal{G}(\mathbf{z}_{\tau},\tau,\mathbf{H},\mathbf{c})$, where $\mathbf{c}$ is an optional text embedding. During inference we sample $\mathbf{z}_{\tau_{\max}}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$, perform about $\tau_{\max}\!\approx\!10$ denoising steps to obtain $\widehat{\mathbf{z}}_{0}$, decode $\widehat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\widehat{\mathbf{z}}_{0})$, and update $\mathbf{H}$ with the last $H$ frames for autoregressive sampling. This latent backbone is reused; our contribution lies in coupling it with 3DGS-specific controllers.
@@
Scene-adapted RL locomotion in 3DGS. Our insight here is that locomotion policies trained in mesh-based synthetic environments Zhao et al. (2023) can be used in 3DGS reconstructions when combined with our scene adaptation. We cast navigation as an MDP whose action space is the latent space of the motion model. The policy outputs a diffusion start-noise $\mathbf{z}^{(\tau_{\max})}_{\text{RL},i}$, which a frozen $\mathcal{G},\mathcal{D}$ map to a short motion clip, ensuring stable rollouts. At step $i$, the agent observes state $s_{i}=(\mathbf{H}_{i},\mathbf{g}_{i},\bm{\eta}_{i},\mathbf{c}_{i})$ where $\mathbf{H}_{i}$ is motion history, $\mathbf{g}_{i}$ a goal cue, $\bm{\eta}_{i}$ a scene cue, and $\mathbf{c}_{i}$ a text embedding. The policy samples $a_{i}\sim\pi_{\theta}(\cdot\mid s_{i})$, interpreted as $\mathbf{z}^{(\tau_{\max})}_{\text{RL},i}$. The resulting clip $\mathbf{X}_{i}$ updates the history $\mathbf{H}_{i+1}$. Rewards $r_{i}=r(s_{i},a_{i},s_{i+1})$ encourage waypoint progress, obstacle avoidance, and kinematic plausibility. Training follows synthetic mesh-based environments as in Zhao et al. (2023), while our contribution is the deployment in 3DGS. For deployment in 3DGS scenes (which lack meshes), we approximate navigation meshes via orthographic projection: (i) compute PCA of Gaussian centers to align a top-down view, (ii) threshold opacities to filter floaters, (iii) render a binary map of obstacles and run A* for pathfinding. The policy consumes an egocentric occupancy grid/walkability map $\mathcal{M}\in\{0,1\}^{N\times N}$ centered on the agent. For each grid cell $u$, we compute its nearest-neighbor distance $d(u)$ to filtered Gaussians and mark $\mathcal{M}(u)=1$ if $d(u)>\tau$, else 0. Despite being approximate, this provides sufficiently reliable local context for navigation in 3DGS scenes. For inference during locomotion, we fix text cue $\mathbf{c}_{i}$ to ”walk”. For further details please see supp mat.
@@
Latent optimization for transitions in 3DGS. Once the agent reaches the vicinity of an action point, control switches from RL to deterministic latent-space optimization for fine-grained actions such as stopping, sitting, or grasping. Following Zhao et al. (2025), we adopt a deterministic DDIM sampler (no step-skipping), which defines a fixed rollout (see supp. mat.) $\mathbf{M}=\mathrm{ROLLOUT}(\mathbf{Z}_{\text{opt}},\mathbf{H}_{\mathrm{seed}},\mathbf{C})$, where $\mathbf{Z}_{\text{opt}}$ is the terminal noise variable, $\mathbf{H}_{\mathrm{seed}}$ the seed history, and $\mathbf{C}$ a fixed text cue (“sit”, “grab”). We optimize $\mathbf{Z}_{\text{opt}}$ by minimizing $\mathcal{L}(\mathbf{Z}_{\text{opt}})=F(\Pi(\mathbf{M}),\mathbf{g_{user}})+\mathrm{Cons}(\mathbf{M})$ with gradient updates $\mathbf{Z}_{\text{opt}}^{(k+1)}=\mathbf{Z}_{\text{opt}}^{(k)}-\eta\nabla_{\mathbf{Z}_{\text{opt}}}\mathcal{L}(\mathbf{Z}_{\text{opt}}^{(k)})$, where $\Pi(\cdot)$ projects the rollout onto task-relevant variables, $F$ measures goal satisfaction, and $\mathrm{Cons}$ adds continuity, collision, and smoothness constraints.
@@
For position-only goals $\mathbf{g_{user}}$ (e.g., sitting or grabbing at a user-provided point), we synthesize a short $f$-frame snippet $\mathbf{M}=(\mathbf{M}_{1},\dots,\mathbf{M}_{f})$ starting from the locomotion end state $\mathbf{M}^{\mathrm{loc}}_{\mathrm{end}}$. In this setting, $F$ corresponds to the reach and stop terms, with $\mathcal{L}_{\mathrm{reach}}=\|\mathbf{x}_{f}(j^{\star})-\mathbf{g}\|_{2}^{2}$ and $\mathcal{L}_{\mathrm{stop}}=\|\mathbf{v}_{f}(j^{\star})\|_{2}^{2}$, while $\mathrm{Cons}$ corresponds to start-continuity $\mathcal{L}_{\mathrm{start}}=\|\mathbf{M}_{1}-\mathbf{M}^{\mathrm{loc}}_{\mathrm{end}}\|_{2}^{2}$, collision $\mathcal{L}_{\mathrm{coll}}=\sum_{b\in\mathcal{B}_{f}}[-\phi(b)]_{+}^{2}$, and smoothness $\mathcal{L}_{\mathrm{smooth}}=\tfrac{1}{f-1}\sum_{t=2}^{f}\|\mathbf{M}_{t}-\mathbf{M}_{t-1}\|_{2}^{2}$. The final objective is therefore $\mathcal{L}=\mathcal{L}_{\mathrm{reach}}+\lambda_{v}\mathcal{L}_{\mathrm{stop}}+\lambda_{\mathrm{start}}\mathcal{L}_{\mathrm{start}}+\lambda_{\mathrm{coll}}\mathcal{L}_{\mathrm{coll}}+\lambda_{s}\mathcal{L}_{\mathrm{smooth}}$. Here $j^{\star}$ is an anchor joint (e.g., pelvis), $\mathbf{x}_{f},\mathbf{v}_{f}$ are its pose and velocity at frame $f$, $\mathcal{B}_{f}$ are sampled SMPL points at frame $f$, and $\phi(\cdot)$ is a differentiable signed-distance proxy to 3DGS Gaussians . After completing the action, the same formulation synthesizes the exit transition (e.g., sit $\to$ walk), after which locomotion resumes.
@@ 3.3 Differentiable Contact Refinement in 3DGS
After animating the reconstructed human Gaussians with synthesized human motion data (Sec. 3.2), we place them into the reconstructed 3DGS scene. A naive composition of posed human Gaussians with scene Gaussians often leads to floor/geometry penetration and inconsistent contacts. We introduce a contact-aware refinement that solves for small, physically meaningful translations of a sparse set of human Gaussians so that contacts are respected and penetrations are reduced. Setup. Let the posed human Gaussians at time $t$ be $\mathcal{G}^{\mathsf{P}}_{t}=\{(\mathbf{x}_{k}^{\mathsf{P}},\bm{\Sigma}_{k}^{\mathsf{P}},\alpha_{k},\mathbf{c}_{k})\}_{k=1}^{N_{H}}$, and the scene Gaussians be $\mathcal{G}^{S}$. Our goal is to refine a subset of the human Gaussians by per-frame translations $\mathbf{T}_{k,t}$ to achieve (i) contact where appropriate and (ii) separation elsewhere.
@@
Figure 2: Qualitiative results: Our refinement yields consistent contacts across diverse scenes and identities (4–48 camera captures).

### v2

@@ 3.1 Gaussian Reconstruction
Step 1: Canonical Gaussian parameterization on SMPL. Given multi-view images of a person performing diverse poses, we learn a mapping from SMPL poses to 3D Gaussians in posed space. Following (li2023animatable), we place Gaussians on the 2D manifold of the SMPL surface by constructing an approximate front–back UV atlas via orthographic projections of the SMPL mesh. Let $\bm{\beta}$ denote SMPL shape and $\bm{\theta}_{t}$ the pose at time $t$. We rasterize pose-conditioned features into a pose map $P_{t}\in\mathbb{R}^{H\times W\times C}$ - denoted using $\mathcal{M}(\bm{\beta},\bm{\theta_{t}})$. A per-identity StyleUNet $f_{\phi}$ predicts a set of canonical human Gaussians anchored on the SMPL surface: $\mathcal{G}^{\mathsf{C}}_{t}\;=\;f_{\phi}(P_{t})\;=\;\big\{(\mathbf{x}_{k}^{\mathsf{C}},\bm{\Sigma}_{k}^{\mathsf{C}},\mathbf{c}_{k},\alpha_{k})\big\}_{k=1}^{N_{H}}.$
@@
Step 3: Pose-to-Gaussian inference (test-time). Following (li2023animatable), we compute the top $K\!\in\![10,20]$ PCA components of training pose maps $\{P_{t}\}$, yielding mean $\bar{P}$ and basis $Q_{K}$. At inference, poses synthesized by our Gaussian-aligned motion module (Sec. 3.2) are first projected to this subspace and then mapped to posed Gaussians: $\tilde{P}_{y}\;=\;\bar{P}+Q_{K}z_{y}\qquad\mathcal{G}^{\mathsf{P}}_{y}\;=\;\mathrm{LBS}_{\bm{\theta}_{y}}\!\big(f_{\phi}(\pi\!\big(\mathcal{M}(\beta,\bm{\theta}_{y})\big))\big).$ Here we use $\pi$ to denote the projection to the subspace operation and $y$ to indicate a test-time pose. For further details please see supplementary materials.
@@ 3.2 Gaussian-Aligned Motion Synthesis
Design overview. We reuse a strong latent motion backbone trained on large scale mocap dataset and add two 3DGS-specific controllers: an RL locomotion policy that navigates between waypoints while avoiding scene Gaussians, and a deterministic latent-space optimizer that executes short, fine-grained actions near targets (e.g., stop, sit, grasp) before returning control to RL. While this explicit decomposition is not typical in existing motion synthesis frameworks, we find it especially effective in 3DGS settings, as this allows us to exploit the fact that in 3DGS scenes much of the raw scene detail can be abstracted to (i) a set of paths for navigation and (ii) action points (e.g., sitting locations, grasping targets provided by an animator) at which specific behaviors are executed - thus allowing for scene-aware motion synthesis without human-scene paired data. Both submodules operate consistently in the latent space of a learned motion model (Zhao:DartControl:2025).
@@
Latent motion backbone. We adopt a latent motion prior, following prior work (Zhao:DartControl:2025) trained on AMASS (BABEL:CVPR:2021; mahmood19amass). Specifically, the model learns a compact motion-primitive space with a transformer VAE trained on mocap data, and places a diffusion prior in this latent space. Given motion history $\mathbf{H}$ and a future motion segment $\mathbf{X}$, the encoder $\mathcal{E}$ outputs a Gaussian posterior $q_{\phi}(\mathbf{z}\mid\mathbf{H},\mathbf{X})=\mathcal{N}(\bm{\mu},\sigma^{2}\mathbf{I})$ with reparameterized sample $\mathbf{z}=\bm{\mu}+\sigma\odot\bm{\varepsilon}$ where $\bm{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The decoder $\mathcal{D}$ reconstructs motion as $\widehat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\mathbf{z})$. On this latent space, a denoiser $\mathcal{G}$ operates with forward process
$q(\mathbf{z}_{\tau}\mid\mathbf{z}_{\tau-1})=\mathcal{N}(\sqrt{1-\beta_{\tau}}\,\mathbf{z}_{\tau-1},\,\beta_{\tau}\mathbf{I})$
and predicts the clean code $\widehat{\mathbf{z}}_{0}=\mathcal{G}(\mathbf{z}_{\tau},\tau,\mathbf{H},\mathbf{c})$, where $\mathbf{c}$ is an optional text embedding. During inference we sample $\mathbf{z}_{\tau_{\max}}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$, perform about $\tau_{\max}\!\approx\!10$ denoising steps to obtain $\widehat{\mathbf{z}}_{0}$, decode $\widehat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\widehat{\mathbf{z}}_{0})$, and update $\mathbf{H}$ with the last $H$ frames for autoregressive sampling. This latent backbone is reused; our contribution lies in coupling it with 3DGS-specific controllers.
@@
Scene-adapted RL locomotion in 3DGS. Our insight here is that locomotion policies trained in mesh-based synthetic environments (zhao2023dimos) can be used in 3DGS reconstructions when combined with our scene adaptation. We cast navigation as an MDP whose action space is the latent space of the motion model. The policy outputs a start-noise $\mathbf{z}^{(\tau_{\max})}_{\text{RL},i}$, which a frozen $\mathcal{G},\mathcal{D}$ map to a short motion clip, ensuring stable rollouts. At step $i$, the agent observes state $s_{i}=(\mathbf{H}_{i},\mathbf{g}_{i},\bm{\eta}_{i},\mathbf{c}_{i})$ where $\mathbf{H}_{i}$ is motion history, $\mathbf{g}_{i}$ a goal cue, $\bm{\eta}_{i}$ a scene cue, and $\mathbf{c}_{i}$ a text embedding. The policy samples $a_{i}\sim\pi_{\theta}(\cdot\mid s_{i})$, interpreted as $\mathbf{z}^{(\tau_{\max})}_{\text{RL},i}$. The resulting clip $\mathbf{X}_{i}$ updates the history $\mathbf{H}_{i+1}$. Rewards $r_{i}=r(s_{i},a_{i},s_{i+1})$ encourage waypoint progress, obstacle avoidance, and kinematic plausibility. Training follows synthetic mesh-based environments as in (zhao2023dimos), while our contribution is the deployment in 3DGS. For deployment in 3DGS scenes (which lack meshes), we approximate navigation regions via orthographic projection: (i) compute PCA of Gaussian centers to align a top-down view, (ii) threshold opacities to filter floaters, (iii) render a binary map of obstacles and run A* for pathfinding. The policy consumes an egocentric occupancy grid/walkability map $\mathcal{M}\in\{0,1\}^{N\times N}$ centered on the agent. For each grid cell $u$, we compute its nearest-neighbor distance $d(u)$ to filtered Gaussians and mark $\mathcal{M}(u)=1$ if $d(u)>\tau$, else 0. Despite being approximate, this provides sufficiently reliable local context for navigation in 3DGS scenes. For inference during locomotion, we fix text cue $\mathbf{c}_{i}$ to ”walk”. For details please see supp mat.
@@
Latent optimization for transitions in 3DGS. Once the agent reaches the vicinity of an action point, control switches from RL to deterministic latent-space optimization for fine-grained actions such as stopping, sitting, or grasping. Following (Zhao:DartControl:2025), we adopt a deterministic DDIM sampler (no step-skipping), which defines a fixed rollout (see supp. mat.) $\mathbf{M}=\mathrm{ROLLOUT}(\mathbf{Z}_{\text{opt}},\mathbf{H}_{\mathrm{seed}},\mathbf{C})$, where $\mathbf{Z}_{\text{opt}}$ is the terminal noise variable, $\mathbf{H}_{\mathrm{seed}}$ the seed history, and $\mathbf{C}$ a fixed text cue (“sit”, “grab”). We optimize $\mathbf{Z}_{\text{opt}}$ by minimizing
$\mathcal{L}(\mathbf{Z}_{\text{opt}})=F(\Pi(\mathbf{M}),\mathbf{g_{user}})+\mathrm{Cons}(\mathbf{M})$
with gradient updates $\mathbf{Z}_{\text{opt}}^{(k+1)}=\mathbf{Z}_{\text{opt}}^{(k)}-\eta\nabla_{\mathbf{Z}_{\text{opt}}}\mathcal{L}(\mathbf{Z}_{\text{opt}}^{(k)})$, where $\Pi(\cdot)$ projects the rollout onto task-relevant variables, $F$ measures goal satisfaction, and $\mathrm{Cons}$ adds continuity, collision, and smoothness constraints.
@@
For position-only goals $\mathbf{g_{user}}$ (e.g., sitting or grabbing at a user-provided point), we synthesize a short $f$-frame snippet $\mathbf{M}=(\mathbf{M}_{1},\dots,\mathbf{M}_{f})$ starting from the locomotion end state $\mathbf{M}^{\mathrm{loc}}_{\mathrm{end}}$. In this setting, $F$ corresponds to the reach and stop terms, with $\mathcal{L}_{\mathrm{reach}}=\|\mathbf{x}_{f}(j^{\star})-\mathbf{g}\|_{2}^{2}$ and $\mathcal{L}_{\mathrm{stop}}=\|\mathbf{v}_{f}(j^{\star})\|_{2}^{2}$, while $\mathrm{Cons}$ corresponds to start-continuity $\mathcal{L}_{\mathrm{start}}=\|\mathbf{M}_{1}-\mathbf{M}^{\mathrm{loc}}_{\mathrm{end}}\|_{2}^{2}$, collision $\mathcal{L}_{\mathrm{coll}}=\sum_{b\in\mathcal{B}_{f}}[-\phi(b)]_{+}^{2}$, and smoothness $\mathcal{L}_{\mathrm{smooth}}=\tfrac{1}{f-1}\sum_{t=2}^{f}\|\mathbf{M}_{t}-\mathbf{M}_{t-1}\|_{2}^{2}$. The final objective is therefore
$\mathcal{L}=\mathcal{L}_{\mathrm{reach}}+\lambda_{v}\mathcal{L}_{\mathrm{stop}}+\lambda_{\mathrm{start}}\mathcal{L}_{\mathrm{start}}+\lambda_{\mathrm{coll}}\mathcal{L}_{\mathrm{coll}}+\lambda_{s}\mathcal{L}_{\mathrm{smooth}}$
Here $j^{\star}$ is an anchor joint (e.g., pelvis), $\mathbf{x}_{f},\mathbf{v}_{f}$ are its pose and velocity at frame $f$, $\mathcal{B}_{f}$ are sampled SMPL points at frame $f$, and $\phi(\cdot)$ is a differentiable signed-distance proxy to 3DGS Gaussians . After completing the action, the same formulation synthesizes the exit transition (e.g., sit $\to$ walk), after which locomotion resumes. Though our method can synthesize actions which mimic “picking up”, or “grabbing” in the scene it cannot lift actual objects from the scene - as we assume that the scene remains static throughout.
Figure 2: With (left) and right (without) refinment of Gaussians
@@ 3.3 Differentiable Contact Refinement in 3DGS
After animating the reconstructed human Gaussians with synthesized human motion data (Sec. 3.2), we place them into the reconstructed 3DGS scene. Here we want to highlight that we never use the SMPL mesh. We use the “SMPL pose” (the joint angles of the 24 SMPL joints or the 54 SMPL-X joints) - to drive the 3DGS avatar. A naive composition of posed human Gaussians with scene Gaussians often leads to floor/geometry penetration and inconsistent contacts. We introduce a contact-aware refinement that solves for small, physically meaningful translations of a sparse set of human Gaussians so that contacts are respected and penetrations are reduced. (Fig. 2)
We first note that mapping SMPL pose to 3D Gaussians is fundamentally different from mapping SMPL pose to 3D SMPL mesh. Mapping SMPL pose to a mesh is just a linear blend skinning operation that maps canonical SMPL mesh vertices to posed SMPL vertices - hence the SMPL pose can usually be easily optimized to maintain contact with the scene. This is not the case with 3D human Gaussians. Mapping SMPL pose to 3D human Gaussians involves a forward pass through the learnt network - and then applying LBS to the output Gaussians. If we were to optimize the SMPL pose for correct contact this would entail getting a gradient through the neural network that maps SMPL pose to 3D Gaussians - which would probably be difficult to converge. Instead we optimize per-Gaussian offsets as the Gaussians (after being output by the network) are already placed close enough to reasonable locations in the 3d scene; thus we can simply optimize per gaussian offsets - with heavy regularization for correct contact. We describe the setup in detail below.
Setup. Let the posed human Gaussians at time $t$ be $\mathcal{G}^{\mathsf{P}}_{t}=\{(\mathbf{x}_{k}^{\mathsf{P}},\bm{\Sigma}_{k}^{\mathsf{P}},\alpha_{k},\mathbf{c}_{k})\}_{k=1}^{N_{H}}$, and the scene Gaussians be $\mathcal{G}^{S}$. Our goal is to refine a subset of the human Gaussians by per-frame translations $\mathbf{T}_{k,t}$ to achieve (i) contact where appropriate and (ii) separation elsewhere.
@@
Figure 3: Qualitative results: Our method generates diverse motions across scenes and identities.